### v2

# SN 2021lwz: Another Exotic Luminous and Fast Evolving Optical Stripped Envelope Supernova ?

Abstract
Context. Current large-scale, high-cadence surveys, such as the Zwicky Transient Facility (ZTF), provide detections of new and rare types of transients and supernovae whose physical origins are not well understood.
Aims. We investigate the nature of SN 2021lwz at a redshift , an overluminous supernova (SN) of absolute magnitude, AB, falling in the lower range of superluminous supernovae (SLSNe) luminosities, and discovered in a faint dwarf galaxy with an absolute magnitude of AB.
Methods. SN 2021lwz is studied using optical spectroscopy, photometry and imaging linear polarimetry obtained during several follow-up campaigns. All the data are used to analyse and model the evolution of the explosion. Comparisons with other SNe of well known or rarer types are investigated.
Results. SN 2021lwz belongs to the rare class of rapidly evolving transients. The bolometric light curve rises in about days to a peak luminosity of about erg/s, at a rate of 0.2 mag day-1 close to the peak. Spectroscopy modelling reveals more similarities with a normal Type Ic-like SN than with a SLSN before peak, showing slightly broadened lines after peak. Light curve modelling shows that the Arnett model of the bolometric light curve using a radioactive source (56 Ni) is not able to reasonably explain the light curve evolution. A magnetar model seems more appropriate, suggesting that the explosion of low ejecta mass () took place in a low mass () dwarf galaxy of specific star-formation rate about ten times larger than typical star-forming galaxies.
Conclusions. In conclusion SN 2021lwz is an uncommon transient showing many similarities with several classes of transients, and with rare transients. It may be an interesting example pointing on how differences in ejecta mass and engine parameters could produce a wide range of engine-driven SESNe.
Key Words.:
supernova: general – supernova: Individual (SN 2021lwz, iPTF 16asu, SN 1998bw, SN 2018gep, SN 2018bgv, SN 2014ft, AT 2020xnd) – techniques: spectroscopy, photometry, polarimetry1 Introduction
The advent of large-scale high-cadence surveys like Catalina Real-Time Transient Survey (CRTS; Drake et al. 2009), the Panoramic Survey Telescope and Rapid Response System (PanSTARRS; Kaiser et al. 2010), the Palomar Transient Factory (PTF) (see Rau et al. 2009), the Zwicky Transient Facility (ZTF; Bellm et al. 2019b; Graham et al. 2019) and the Asteroid Terrestrial- impact Last Alert System (ATLAS; Tonry et al. 2018a), developed over the last two decades, have opened the possibility to detect new types of transients. They have been useful to start populating the parameter space of the rising time, , defined as the time between the explosion and peak magnitude, versus absolute magnitude at peak, , with new populations of objects. Among those, superluminous supernovae (SLSNe) are explosions expected to originate from core collapsing stars. Such explosions may be forming a newborn rotating neutron star (Ostriker and Gunn 1971), or a magnetar (Kasen and Bildsten 2010; Woosley 2010). Interactions of the explosion products with the circumstellar medium (CSM) coming from previous progenitor mass loss is an additional mechanism often considered to interpret the diversity of light curves of H-poor SLSNe (e.g. Hosseinzadeh et al. 2022). A SLSN has often been considered as a type of stellar explosion that exhibits a peak luminosity ten or more times higher than that of a typical SN. This definition is quite general. In practice H-poor SLSNe are detected with absolute r-band magnitudes lying between and magnitudes (Gomez et al. 2024). The light curves of the largest sample from ZTF (78 SLSNe) evolve slowly with a mean rest-frame rise time of 41.9 17.8 days (see Chen et al. 2023a, b), while Gomez et al. (2024) derive a mean rise time of 27 days on a larger sample. In both distributions low luminosity ( mag) extremely fast rising ( 10 days) transients are rare. From the ZTF sample, with a rising time of 10 days, and an absolute magnitude peaking close to , SN 2018bgv was the fastest SLSN observed by ZTF during the first months of the survey (Lunnan et al. 2020).
The formation of stars as single objects, or in binary, and even triple stars systems, which is followed by their evolution up to a final stage (e.g. core collapse or disruptive events) suggest the existence of a complex diversity of star evolution channels. Stars evolving in binary systems are subject to strong interactions with their environment (e.g. binary star companion, black hole proximity, winds, shock waves) potentially leading to new types of transients. As such, a variety of new types of transients has recently been discovered among which Luminous Fast Blue Optical Transients (LFBOTs like AT 2018cow (the Cow) and AT 2020xnd (the Camel); see for example Perley et al. 2021b; Ho et al. 2023, and references therein), Ultra-Stripped Supernovae (USSNe; like SN 2014ft (or iPTF 14gqr) and SN 2019ehk, see for example De et al. 2018; Nakaoka et al. 2021, and references therein). A more general terminology encompassing these events is Rapidly Evolving Transients (RETs, e.g. Drout et al. 2014; Arcavi et al. 2016; Tanaka et al. 2016; Pursiainen et al. 2018). Ideally, each of these types would be filling a given area in the phase space ( , ) but, in practice distinct populations sometimes fill similar areas leading to a degeneracy in this parameter space, and more data-driven information is required. This happens for example with well studied populations like Type Ia and Type II SNe, and with rarer populations like Type Ib, Ic and broad line Ic (or Ic-BL) SNe (see e.g. Figure 1 in Whitesides et al. 2017). Such a plot is presented in Figure 1 where the position of SN 2021lwz (see Section3.4) and several other objects of various types are displayed.
The empirical method used to study such new populations of transients, and to try to understand the nature of their progenitors and their environments, is therefore to gather as much data as possible. This includes getting multi-band, multi-messenger, high cadence photometry allowing to estimate the full energy budget spatially deployed by the explosion over time (e.g. Prentice et al. 2016; Kornilov et al. 2023). This requires to get high cadence spectroscopy to study the chemical evolution of the photosphere (e.g. Quimby et al. 2011; Könyves-Tóth 2022; Könyves-Tóth and Seli 2023). When possible this means to get polarimetry data to study the time-evolution of the geometry of the systems closely connected to the explosion (e.g. inner core, jets, see Nagao et al. 2024; Soker 2024) or products resulting from the explosion (e.g. photosphere evolution, ejecta, see Inserra et al. 2016; Leloudas et al. 2017). Obviously gathering such data is a complex task even more complex for fast-rising events, and the lack of data obtained on some specific objects often prevents us from getting the full picture in one go. Nevertheless, gathering data and increasing the sample on several members of the same class of objects, retrospectively helps to better understand the said class. Such a progress is done in parallel to theoretical and numerical modelling developments (e.g. Khatami and Kasen 2024; Omand and Jerkstrand 2023), which are crucial to understand and interpret the data, and ultimately to put constraints on, or to identify the channel formation of such new or peculiar transients or SNe. In this context of luminous transients SN 2021lwz is a particularly intriguing object deserving a dedicated in depth study.
SN 2021lwz (also known as ZTF21abaiono, PS21fcj, ATLAS21oii and Gaia21cvf), at R.A.09h44mn47.390s, Dec, was discovered by Fremling (2021) on May 10, 2021 from ZTF 111Zwicky Transient Facility, https://www.ztf.caltech.edu. public alerts (Patterson et al. 2019). The discovery magnitude obtained with the ZTF-cam mounted on the Palomar 1.2 meter Oschin was of g18.76 mag (AB system). The transient was classified on May 15, 2021 as a SLSN-I at a redshift z by Perley et al. (2021a) from the analysis of a Keck1 Low Resolution Imaging Spectrometer (LRIS, Oke et al. 1995) 222LRIS, https://www2.keck.hawaii.edu/inst/lris/lrishome.html spectrum. For that reason SN 2021lwz entered in our sample of H-poor SLSNe for follow-up. Subsequently, though, we realized that some of the pre-peak spectra are not similar to other typical H-poor SLSN spectra at similar phases, and we therefore decided to investigate this source in more detail. Indeed, SN 2021lwz is also part of the sample of rare Luminous SNe discussed by Gomez et al. (2022). This type of transient appearing to form a continuum between type Ic SNe and SLSNe is expected to make up of all observed Core Collapse SNe (CCSNe) (Gomez et al. 2022). Additionally, SN 2021lwz is a fast-rising SLSN ranging in the low absolute magnitude range, as do most of the He-rich SLSNe. This type is rare but a sample of seven Helium-Rich SLSNe was reported by Yan et al. (2020). In this sample, the spectroscopically well sampled SLSN SN 2019hge (ZTF 19aawfbtg) was shown to be of similar spectral type as PTF 10hgi, the only known He-rich SLSN at the time. Both were used to identify 6 more He-rich SLSNe thus increasing the sample of such known events to a total of 8 members.
In this work we present an intensive study of SN 2021lwz. All the data obtained on SN 2021lwz and its host galaxy are presented in section 2. The data analysis of SN 2021lwz provided in section 3 is followed by detailed modelling of the SN and its host galaxy in section 4. In order to better understand and contextualize the nature of SN 2021lwz, comparisons with some SLSNe and other typical SNe and exotic transients are discussed in section 5. Our conclusions are given in section 6.
Along this work, the photometric measurements are reported in the AB system and the uncertainties are provided at 1 confidence. We use the Planck 2018 Flat CDM cosmology model ( kms) (Planck Collaboration et al. 2020)
2 Data
In this section, we present data and information about the host galaxy. We also present the photometry, spectroscopy and polarimetry data obtained at several epochs on SN 2021lwz. Most of these observations were coordinated through the ZTF Fritz Marshal system (van der Walt et al. 2019; Coughlin et al. 2023). The reference epoch of maximum light is defined as MJD days with respect to g-band, as from a visual inspection. The extinction contribution from our Galaxy in the line-of-sight to SN 2021lwz and its host is and (Schlafly and Finkbeiner 2011).
2.1 Host photometry
The highly accurate spectroscopic redshift z of SN 2021lwz was determined from narrow H-alpha host-galaxy emission and probable [OIII] emission at a redshift consistent with the less accurate redshift obtained with the SNID analysis (Perley et al. 2021a). The host galaxy at R.A.09h44mn47.397s, Dec is detected in archival Sloan Digitized Sky Survey (SDSS) 333http://skyserver.sdss.org/dr17/VisualTools/explore/summary and Legacy Survey (LS) 444Legacy Survey, https://www.legacysurvey.org/ images. It can be seen as the blue feature marked by the yellow circle in the LS image plotted in Figure 2.
We retrieved science-ready co-added images from the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS, PS1) DR1 (Chambers et al. 2016) and the Sloan Digital Sky Survey data release 9 (SDSS DR 9; Ahn et al. 2012). In addition, the field was observed with the 3.58-m Canada France Hawaii Telescope telescope equipped with MegaPrime in u and r band between 2016 and 2019. We retrieved the level-2 data from the CADC and combined them using Swarp version 2.38.0 (Bertin 2010) to create a co-added image. We measure the brightness in all images using an aperture photometry tool presented in Schulze et al. (2018) that is based on Source Extractor (Bertin and Arnouts 2010). The measurements in these different bands are given in Table 4 (see Appendix A).
One survey observed that part of the sky in the near-infrared (NIR) with the Wide Field Camera (WFCAM) mounted on the United Kingdom Infrared Telescope 555UKIRT, https://about.ifa.hawaii.edu/ukirt/. The Large Area Scale (LAS) survey footprint of the UKIDSS Data Release 9 666https://cdsarc.cds.unistra.fr/viz-bin/cat/II/319 includes the coordinates position of SN 2021lwz. This survey whose limit magnitudes are all lower than 20.2 mag in the Y, J, H and K bands does not show any detection of the host of SN 2021lwz (see Lawrence et al. 2007), in agreement with our results from the modelling of the host (see Section 4.4).
2.2 Spectroscopy
In addition to the Keck1 LRIS public spectrum used to classify the transient, we acquired spectra with the Spectral Energy Distribution Machine (SEDM; Blagorodnova et al. 2018) on the Palomar Observatory 60 inch telescope (P60), the LRIS on the 10 m Keck I telescope and the Andalucia Faint Object Spectrograph and Camera (ALFOSC)777http://www.not.iac.es/instruments/alfosc/ on the 2.56 m Nordic Optical Telescope (NOT). SEDM spectra were reduced according to Rigault et al. (2019); Kim et al. (2022). The LRIS spectrum was reduced using LPipe as described by Perley (2019). ALFOSC spectra were reduced using PyNOT888https://github.com/jkrogager/PyNOT. A spectrophotometric standard star was used to calibrate each spectrum. A summary of the spectra obtained on SN 2021lwz is given in Table 5. In Figure 3 we show the spectral evolution of SN 2021lwz. The spectra were taken between rest frame days and days relative to the g-band maximum.
The brightness and low-redshift of SN 2021lwz motivated observations in the NIR. One spectrum was obtained at days, rest frame, with the Near-Infrared Echellette Spectrometer (NIRES Wilson et al. 2004), a prism cross-dispersed near-infrared spectrograph mounted on the Keck II telescope. The spectrograph has a power resolution of R . It simultaneously covers the bands over the wavelength range of 0.8 – 2.4 m. The spectra were reduced with spextool (Cushing et al. 2004) and Xtellcor (Vacca et al. 2003), then combined together. The final spectrum is shown in Figure 4.
2.3 SN photometry
Photometry of SN 2021lwz was obtained in several filters with the Swift satellite (Gehrels et al. 2004), from the ZTF survey (Bellm et al. 2019b), and with the Liverpool Telescope (LT) (Steele et al. 2004). Public photometry was retrieved from the Asteroid Terrestrial-impact Last Alert System (ATLAS) 999https://atlas.fallingstar.com (Tonry et al. 2018b) public server (as explained in Section 2.3.4 ) Details about the data are given in the following sections. The photometry is shown in Figure 5. Table 6 summarises all these measurements.
2.3.1 Swift UVOT data
We observed the field with the 30 cm Ultraviolet/Optical Telescope (UVOT; Roming et al. 2005) aboard the The Neil Gehrels Swift Observatory, hereafter Swift satellite (Gehrels et al. 2004) between 2.6 and 20.9 days before and after maximum light (observer-frame) in 6 different filters (, , , , , ) from 2086 to 5411 Å. We retrieved the science-ready data from the Swift archive 101010https://www.swift.ac.uk/swift_portal. We co-added all sky exposures for a given epoch and filter to boost the S/N using uvotimsum in HEAsoft111111https://heasarc.gsfc.nasa.gov/docs/software/heasoft/ version 6.26.1. Afterwards, we measured the brightness of SN 2021lwz with the Swift tool uvotsource. The source aperture had a radius of while the background region had a significantly larger radius. The photometry was calibrated with the latest calibration files from September 2020 and converted to the AB system using Breeveld et al. (2011).
2.3.2 ZTF data
Data were acquired with the Zwicky Transient Facility (ZTF) that uses the Samuel Oschin 48-inch (1.22 m) Schmidt telescope at Palomar Observatory on Mount Palomar (USA). It is equipped with a 47-square-degree camera (Dekany et al. 2020) and monitors the entire northern hemisphere every 2–3 days in the and bands to a depth of mag (; Bellm et al. 2019b; Graham et al. 2019) as part of the public ZTF Northern Sky Survey (Bellm et al. 2019a).
We retrieved the host-subtracted photometry via the ZTF forced-photometry service (Masci et al. 2023) hosted by NASA/IPAC Infrared Science Archive (IRSA) at IPAC.
2.3.3 Liverpool Telescope data
Observations were taken with the IO:O optical imager on the LT in the Sloan Digital Sky Survey (SDSS) , , , and filters. Reduced images were downloaded from the LT archive and processed with custom image-subtraction and analysis software (K. Hinds and K. Taggart et al., in prep.). Image stacking and alignment is performed using SWarp (Bertin et al. 2002) where required. Image subtraction was performed using a pre-explosion reference image in the appropriate filter from the Panoramic Survey Telescope and Rapid Response System 1 (Pan-STARRS1) or SDSS. The photometry is measured using PSF fitting methodology relative to Pan-STARRS1 or SDSS standards and is based on techniques in Fremling et al. (2016).
2.3.4 ATLAS data
ATLAS forced photometry data were retrieved from the ATLAS public server 121212https://fallingstar-data.com/forcedphot/ (Tonry et al. 2018b; Smith et al. 2020; Shingles et al. 2021). They were clipped and binned to one day using the publicly available code plotatlasfp.py 131313https://gist.github.com/thespacedoctor/86777fa5a9567b7939e8d84fd8cf6a76.
2.4 SN polarimetry
The observations log of the imaging polarimetry obtained with ALFOSC on the Nordic Optical Telescope (NOT) is given in Table 7. Linear polarimetry was obtained at +1.0 and +19.0 days, rest-frame. The polarimetry was obtained using a half wave plate in the FAPOL unit and a calcite plate mounted in the aperture wheel. The calcite plate provides the simultaneous measurement of the ordinary and the extraordinary components of two orthogonal polarized beams (see Figure 18). The half wave plate can be rotated in steps of 22.5∘ from 0∘ to . As a standard, we used 4 angle positions (, and ) to sample the Stokes parameters space. The photometry of the ordinary and extra-ordinary beams was done using aperture photometry of size 2 to 3 times the Full-Width at Half-Maximum (FWHM) of point sources in the images. For multiple sequences of 4 Half-Wave Plate (HWP) angles the polarization is obtained by summing-up the fluxes from the ordinary and extra-ordinary beams to minimize uncertainty propagation.
3 Data analysis
In this section we derive physical parameters associated to the explosion using an analytical approach.
3.1 Spectral analysis using template fitting
In order to get a first understanding of the spectral evolution of SN 2021lwz, we ran all the optical spectra through the SuperNova IDentifier (SNID) (Blondin and Tonry 2007, and references therein). We first performed a general classification by running the code over our upgraded template database which, in addition to the standard Types (Ia, Ib, Ic, II) and their subtypes, contains SLSN-I and SLSN-II templates including those of PTF 10hgi (19 spectra with epochs lying between -30 and +312 days with respect to peak in the SN rest-frame) and SN 2019hge (2 spectra at epochs 0 and +48 days, respectively) which are used as He-rich SLSN-I templates in Yan et al. (2020) (see Section 1).
The analysis provided by SNID shows that none of the spectra of SN 2021lwz obtained before peak (see Table 5 and Figure 3) can be matched properly without decreasing some of the values of the default SNID parameters aimed to ensure a statistically robust diagnostic. Fit attempts using only He-rich SLSNe PTF 10hgi or SN 2019hge templates were conducted on the pre-peak spectra without producing any good match. Using the full database, a relatively poor fit is obtained on the -3.4 days Keck spectrum with the Type Ic-SL (Ic-superluminous) PTF10cwr, whose absorption features are smoothed compared to other SL templates. At maximum light the Keck spectrum at -0.5 days show similarities with the fast-evolving transient SN 2018bgv. These fits are shown in Figure 19. Accurate fits of the ALFOSC spectra of SN 2021wz obtained post-peak are obtained without degrading the by-default SNID parameters. They suggest that the photosphere of SN 2021lwz evolved as a Ic-BL looking-like Type, in agreement with the first statement provided in the TNS classification report.
At early times, the spectra look (by eye) like they could be consistent with SLSNe, but with higher velocities than we are used to seeing. This could explain why the fits with SLSNe favour lower redshifts. This may also happen because SNID tries to adjust the W-feature often attributed to O ii, but it is known that different lines of other elements like Fe ii are expected in this wavelength range. This will be discussed further in our analysis. The fits obtained with SLSN-I SN 2016wi and with the Type Ic-SL SN LSQ12dlf, shown in Figure 19; middle, point to more accurate redshifts consistency. Using a slightly lower value of 4 for the rlap factor, it is possible to produce lower quality fits with late time spectra of the peculiar Type Ic SN 1998bw (bottom) whose NIR spectral properties will be discussed and compared with those of SN 2021lwz NIR spectrum in Section 3.2.
3.2 NIRES spectrum
The NIR spectrum of SN 2021lwz, obtained +12.6 days, rest frame, after peak is shown in Figure 4. It provides valuable information for comparison with several types of supernovae observed in the NIR. The spectrum is smooth and relatively featureless with an absorption feature centred at m (see gray vertical line with exclamation mark). If that feature was to be produced by He i, m, it’s velocity would be km s-1, i.e. about the same order as the ejecta velocity derived a few days before maximum light, and about two times larger than the photosphere velocities at late phase ( km s-1), as derived from the optical spectra analysis discussed in the following sections. At this late phase velocity, any He i, m absorption feature would be encountered at a wavelength position centred close to 1.0559 m as indicated by the black vertical line. For comparison, the high velocity km s-1, He i, m absorption feature possibly detected in the peculiar Type Ic SN 1998bw (Patat et al. 2001) is shown by the blue vertical line. Similar comparisons can be done with the He-Rich and He-Poor template spectra at phase +13 days (d) of Stripped-Envelope SNe (SESNe) derived from a principal component analysis of Types IIb, Ib, Ic and Ic-BL by Shahbandeh et al. (2022), and with the days He-Rich spectrum of SLSN-I SN 2019hge (Yan et al. 2020). At higher wavelengths, these 3 spectra show spectral features which are not seen in the spectrum of SN 2021lwz, whose smooth features look more similar to those seen in the spectrum of SN 1998bw.
Additional unidentified absorption or emission features are seen in the spectrum of SN 2021lwz in the wavelength range m, which are not seen in the SESNe template spectra. In this wavelength range, some features are seen in the spectra of SN 2008ax and SN 1998bw in sky-line contaminated regions of the spectra making any direct comparison difficult. At higher wavelength, the SN 2021lwz spectrum does not show any hint of an absorption feature that could be produced by He i, m.
To assess if SN 2021lwz could be a Type I SLSN, we also made comparisons with spectra of SN 2015bn at phases -9 days and +34 days (Nicholl et al. 2016b), of Gaia16apd at phase -2.5 days (Yan et al. 2017, and references therein) and of SN 2012il at phase (Inserra et al. 2013). None of these spectra are at a phase similar to the NIR spectrum of SN 2021lwz, and the wavelength coverage is sometimes limited to make meaningful comparisons, but all in all, they are all looking different from the spectrum of SN 2021lwz.
3.3 P-Cygni velocity estimation
We use the Fritz data science platform to cycle through all of the available lines and find that Fe ii is the best one that matches the broad feature near 4515 Å seen in some of the spectra at a redshift z=0.065.
With this respect, we attribute the prominent P-Cygni-like profiles in the Keck I spectra at and days to be mainly due to Fe ii at 4515 Å . They are marked with black in Figure 3. We estimate the ejecta velocity according to the longitudinal relativistic Doppler shift. We use the scipy.optimize.curve_fit algorithm to fit an equation consisting of a straight line component to match the continuum, and a Gaussian component to match the absorption feature to P-Cygni profile data as shown in Figure 6. The resulting covariance matrices and the corresponding parameter fit values of Å and Å are then used to estimate the absorption line velocity of km s-1 at days and km s-1 at days, respectively.
3.4 Light curves analysis
The absolute magnitude is calculated as
where represents the distance modulus (which is equal to 37.40 for a redshift of 0.065); is the correction factor for the difference in filter bandpass between the observer and the rest frame; A is the extinction caused by the Milky Way; and A is the extinction caused by the host galaxy.
We use only the first term of the K-correction as described by Hogg et al. (2002) as shown here.
The Fitzpatrick (1999) extinction model is used to calculate the MW dust extinction based on the parameters R and E(B-V) mag. We assume that the extinction from the host galaxy is negligible since no absorption lines like for example, Na I Doublet absorption lines, are visible. The resulting absolute magnitude (rest frame) light curves from LT, UVOT, ATLAS and ZTF are plotted in Figure 5.
The rising rate is calculated using the g-band data. At -7.1 days the g-band absolute magnitude is -18.7 mags. At zero days the g-band absolute magnitude is -17.2 mags. The delta mag on the rise close to the peak is therefore 1.5 mags / 7.1 days = 0.2 mags per day.
3.5 Photospheric radius and temperature evolution
We estimate the photospheric temperature and radius evolution at the phases where full sets of UVOT measurements are available, and interpolate ZTF, LT and ATLAS-o light curve measurements using Gaussian Process (GP) regression that utilizes the Python package GEORGE (Ambikasaran et al. 2015) with a Matern 3/2 kernel. We then use the scipy.optimize.minimize141414https://docs.scipy.org/doc/scipy/reference/generated/scipy.optimize.minimize.html algorithm to fit a Planck function (blackbody) to the datasets.
We find that the three UVOT UV filters are consistently incompatible with the best-fit blackbody. We attribute the problem to line blanketing as seen in other SLSNe (see e.g. Yan et al. 2017). We therefore exclude the three UV filter measurements from UVOT before rerunning the blackbody fits to create the photospheric temperature and radius evolutions.
The temperature evolution and the radius evolution of SN 2021lwz are compared to a sample including those of SN 2020qlb (West et al. 2023) and those of the 31 SLSNe-I provided in the online version of Table A67 of Chen et al. (2023a) in Figure 7 and in Figure 8, respectively.
3.6 Ejecta velocity estimation
The time derivative of the best fit polynomial of the SN 2021lwz radius evolution gives rough estimates of the photospheric velocity at relevant phases. At epochs just after explosion the photospheric velocity should approximate the ejecta velocity. At later epochs, P-Cygni absorption line features estimate velocities outside the photosphere. In Figure 9 we compare the photospheric velocity evolution to the P-Cygni velocity estimate from Sec. 3.3.
3.7 Bolometric light curve estimation
At the eight epochs where UVOT data is available we calculate the spectral luminosity (Lλ) for each UVOT filter measurement and GP interpolate filter value from the other filters. Short of the lowest filter value we create Lλ estimates based on the best fit blackbody to the UVOT-UV, UVOT-U and LT-u filter measurements. Long of the highest filter value we create Lλ estimates based on the best fit blackbody to the optical filter measurements. These extrapolation methods are also used in SuperBol software (Nicholl 2018). We then integrate over all wavelengths to create bolometric luminosity estimates, at each of the eight epochs, which form the bolometric light curve. These results are shown, modelled and discussed in section 4.2.
3.8 Photosphere geometry
The polarization degrees measured on the unpolarised star, BD52913, on the polarized calibration star, HD 127769, on the field star STAR 1 (shown in Figure 18), and on SN 2021lwz are all displayed in Table 8. The instrumental polarization () degree is of order of at each epoch. This component, which is retrieved in the parameter space of the Stokes and parameters, is removed from the components of the other sources ( correction). We use the polarized star HD 127769 to calculate the zero polarization angle () which is used to rotate the Stokes parameters from the ALFOSC FAPOL instrument reference frame to the sky reference frame in Equatorial coordinates. The polarization angles are counted positively from North to East. We use the polarimetry available on STAR 1, in the field of view of FAPOL (see Figure 18), to get an estimate of the polarized component of the Galaxy in the vicinity of SN 2021lwz (see values shown in bold in Table 8). These estimates are of the same order as their uncertainties suggesting a low level of the interstellar polarization () in this region of the sky. This result is consistent with the polarization degrees of the two stars HD 82885 () and HD 82635 () retrieved from the Heiles (2000) agglomeration catalog. These two stars are at distances lower than two degrees on the plane-of-sky from the central coordinate position of SN 2021lwz. The levels of polarization obtained on SN 2021lwz after correction using STAR 1 Stoke and parameters, are displayed in the last column of Table 8. The final values shown in bold are obtained after bias correction following the equation given in Wang et al. (1997):
where is the Heaviside function, is the observed polarization and is the 1 error. The final measurements, at +1 day rest-frame, and at +19 days rest-frame, are consistent with null-polarization. They suggest that the geometry of the photosphere of SN 2021lwz which is almost spherical, as from our observer point of view, didn’t change with time. Similarly, there is no hint that the SN’s photosphere interacted with the circumstellar medium (CSM).
4 Data modelling
4.1 Optical spectrum evolution and chemical composition
To study the spectroscopic evolution of SN 2021lwz, we model the available photospheric phase spectra utilizing the code named SYN++ (Thomas et al. 2011). With this code, it is possible to fit some global parameters that refer to the overall model spectrum, such as the temperature () and the velocity () at the photosphere. The lines of the individual ions can be fitted via local parameters, that code the optical depth (), the velocity width of the line forming region ( and ), the scale height parameter of the optical depth above the photosphere (aux) and the excitation temperature (), assuming Local Thermodynamic Equilibrium.
For SN 2021lwz, we have spectra available at 12 different epochs from -5.4 days to +29.1 days phase, rest-frame, relative to the moment of the maximum light. Table 1 presents the date of observation of the modelled high-resolution spectra together with their phases, and the and values of their best-fit models obtained using SYN++. The possibilities, constraints and uncertainties of the usage of SYN++ and its parameters are discussed in details in Könyves-Tóth (2022).
| Date | MJD | Phase | ||
| (days) | (K) | (km s-1) | ||
| 2021-05-13 | 59347 | -3.4 | 20000 | 15000 |
| 2021-05-30 | 59364 | 13.2 | 8000 | 7500 |
| 2021-06-16 | 59381 | 29.1 | 8000 | 7500 |
Figure 10 shows the spectrum modelling of SN 2021lwz. In the top panel, the modelling of the -3.4d phase spectrum of SN 2021lwz can be seen. The best-fit and values can be read out from Table 1. The identified elements are C ii, O i, Na i, Mg ii, Si ii, Ca ii, Ti ii and Fe ii. Regarding this, SN 2021lwz is more similar to a normal Type Ic supernova rather than a SLSN. The pre-maximum spectrum of 2/3 of the SLSNe-I usually contains a W-shaped absorption blend between 4000 and 5000 Å, while the remaining 1/3 fraction of the SLSNe-I show similar features to SN 2015bn (see Könyves-Tóth 2022). However, the -3.4d phase spectrum of SN 2021lwz does not show resemblance to any of these SLSN-I subtypes.
The bottom panel of Figure 10 shows the modelling of the +13.2d and the +29.1d phase spectra, which are quite similar to each other. By these phases, both the photospheric velocity and temperature decreased significantly ( and K), and besides Fe ii, Ca ii, Si ii, Mg ii, O i and C ii, He i was included in the best-fit model.
In Table 9 in the Appendix, the best-fit local parameter values obtained in SYN++ are collected. The SYN++ modelling analysis suggests SN 2021lwz is a intermediate Type Ic/Ic-BL transient rather than a SLSN-I.
To conclude, we point out that any difference between and can be explained with the presence of Non Local Thermodynamic Equilibrium (NLTE) effects. In the case of LTE, for each ion, however, in supernovae, NLTE effects usually play a role in the spectrum formation. Although SYN++ is not a NLTE code, like TARDIS (Kerzendorf and Sim 2014), which can appropriately take into account every difference from LTE, choosing a indicates that there is some divergence from the LTE. If we change the of a given ion in SYN++, the line ratios and strengths become different. We chose the best-fit model accordingly.
4.2 Bolometric light curve modelling
We compare semi-analytic models to the bolometric light curve to determine the most probable power source of the SN. The radioactive source and magnetar source models used by West et al. (2023, Sect. 7) (Arnett 1982; Kasen and Bildsten 2010; Woosley 2010) are employed using the ejecta velocity estimate from Section 3.6. Our main conclusion is that the light curve can’t be fit with . The details of this analysis are given in Appendix G
There are no signs of narrow lines in the spectra indicating interaction, but shock breakout from interaction with extended material (Nakar and Sari 2010; Nakar and Piro 2014; Piro 2015) has been used to explain the light curve of the similar transient SN 2018gep (Ho et al. 2019; Pritchard et al. 2021). We do not explicitly use this model, but note that similar parameters (, 1052 erg, , and 1014 cm (Ho et al. 2019)) could likely also reproduce the light curve of SN 2021lwz.
4.3 Multi-band light curve modelling
We perform inference on the multi-band photometry using the Redback (Sarin et al. 2024) magnetar-driven supernova model including ejecta acceleration and non-vacuum dipole spin-down (Omand and Sarin 2024). We use the PyMultiNest sampler (Buchner et al. 2014) implemented in bilby (Ashton et al. 2019; Romero-Shaw et al. 2020), and sample in flux density with a Gaussian likelihood. The uncertainty presented is only the statistical uncertainty in the fits, and does not include systematic uncertainty inherent in the simplified one-zone Redback model.
We sample the explosion time with a uniform prior of up to 100 days before the first observation, an extinction term with a uniform prior between 0 and 2, and a white noise term. We constrain the total rotational energy of the magnetar to be erg. We fixed the photospheric plateau temperature to a value of 8000 K to keep consistency with the value found in Section 4.1. The physical parameters and priors used are listed in Table 2. Uniform priors are denoted by U and log-uniform priors are denoted by L.
The fitted light curve is shown in Figure 11 and the posterior for the parameters in Figure 21 (see Appendix H). The model fits the peak of the light curve well in all bands but underestimates the flux in the UV bands from around 14 days post-peak. We find an initial luminosity erg s-1 and spin-down timescale one day, and an ejecta mass . These values are similar to that of the luminous fast-blue optical transient (LFBOT) AT 2020xnd (i.e. the ZTF source ZTF20acigmel; see e.g. Perley et al. 2021b), the ultra-stripped supernova (USSN) SN 2014ft (see Table 2 in Omand and Sarin 2024), and iPTF16asu (see Table 2; and Whitesides et al. 2017; Wang et al. 2019; Wang and Gan 2022), including a similar higher magnetar rotational energy of erg, but a lower explosion energy of erg. We find the braking index to be 4.5 and rule out vacuum dipole () with 95% confidence, similar to SN 2015bn and SN 2007ru (Omand and Sarin 2024).
| Parameter Symbol | Definition | Prior | Best Fit Value | Best Fit Value | Units |
| SN 2021lwz | iPTF16asu | Units | |||
| Initial Spin-Down Luminosity | L[40.0,50.0] | 45.45 | erg s-1 | ||
| Spin-Down Timescale | L[2.0,8.0] | 4.64 | s | ||
| Magnetar Braking Index | U[1.5,10] | 8.41 | |||
| Ejecta Mass | L[0.1,100] | 0.17 | |||
| Supernova Explosion Energy | L[0.01,2] | 0.59 | 1051 erg | ||
| Ejecta Opacity | U[0.05,0.20] | 0.06 | cm2 g-1 | ||
| Ejecta Gamma-Ray Opacity | L[-4,4] | 0.27 | cm2 g-1 |
The distribution of ejecta velocities, peak photospheric temperatures, and several timescales is shown in Figure 12. The final ejecta scale velocity is around 11 000 km s-1, which is lower than the inferred photospheric velocity of 15 000 km s-1; this implies that most of the material is travelling at much lower velocity than the photosphere at early times, although by peak the photosphere has receded into the bulk of the material. The photospheric temperature at peak is around 24 000 K, slightly higher than the best-fit model temperature of 20 000 K inferred by the spectral analysis, although not significantly higher. The diffusion time of the supernova,
is very short, only 5 days, due to the low ejecta mass. The bolometric peak timescale is inferred to be 7 days and the -band peak timescale to be 10 days; the discrepancy between these values is consistent with the high photospheric temperature and rapid cooling around peak. From the intrinsic bolometric luminosity we derive days, and days, rest frame. The supernova is expected to enter the nebular phase on the timescale of (Nicholl et al. 2017)
which is 27 days post-explosion, or only 17 days post -band peak. This could imply that the final two observed spectra, at +26d and +29d, are taken during the early nebular phase, even though the spectrum shows no strong evolution between +8 and +29 days. This could also account for the discrepancy between the model and observed photometric points in the late-phase. However, it seems more likely that this simplistic estimate of the nebular timescale is not accurate in this circumstance, which could be because of small scale asymmetries or the density profile of the ejecta. Since Equations 49 are all derived using a uniform density profile, they could all underestimate the different timescales if the ejecta has a steeper profile (e.g. Suzuki and Maeda 2017). Due to the large inferred gamma-ray opacity, the gamma-ray leakage timescale (Wang et al. 2015),
is also very long, with a lower limit of 50 days. This means that a significant amount of non-thermal emission likely does not escape from the supernova over the timescale that we observed it. The inferred lower limit on the gamma-ray opacity is 0.2 cm2 g-1, which is higher than the mean value of 0.06 cm2 g-1 for SLSNe-I (Gomez et al. 2024), although not a significant outlier. The high gamma-ray opacity here may be inconsistent with a pulsar wind nebula (PWN) with low magnetization (Vurm and Metzger 2021), which would cause the non-thermal emission to be Comptonized and mostly emitted as hard x-rays and gamma-rays. The likely source of opacity for the PWN is photoelectric absorption, which dominates in the soft x-ray regime (see Figure 4 from Vurm and Metzger (2021)). This means that the PWN is likely synchrotron-dominated, and may be a strong radio source at longer timescales (e.g. Omand et al. 2018, 2025a).
These parameters differ significantly from Gomez et al. (2022), and Gomez et al. (2024), who used the Modular Open Source Fitter for Transients code (MOSFiT; Guillochon et al. 2018) to fit several luminous supernovae and SLSNe, including SN 2021lwz, using a pure magnetar and magnetar nickel model, respectively. Gomez et al. (2022) infers a higher ejecta mass of 0.4 and a much lower photospheric velocity of 3000 km s-1. Gomez et al. (2024) infers much higher ejecta and 56Ni masses of 3 M and 0.8 M, respectively, while inferring a much less luminous pulsar of 1040 erg s-1, meaning the luminosity of the supernova is mostly powered by 56Ni in that scenario, in contrast to what we find in Section 4.2. While the values from Gomez et al. (2022) give a similar diffusion time to our inferred parameters, there is a strong tension between their derived photospheric velocity and the value we infer from the spectrum. They infer a lower value of erg s-1 for the initial pulsar luminosity but a spin-down timescale of 15 days. The difference in spin-down timescale, as well as their assumption of vacuum-dipole spin down, means the magnetar rotational energy is similar to our inferred parameters. Any discrepancies likely result from the assumption of constant velocity and decoupling of the magnetar spin-down and ejecta dynamics in the mosfit model, which has been shown via radiation hydrodynamics simulations to be unjustified for high-luminosity magnetars (Suzuki and Maeda 2021).
4.4 Host SED modelling
Using the SDSS photometry magnitude (see Table 4) and after taking into account the contribution of the Milky-Way reddening, at a redshift , the absolute magnitude of mag of the host galaxy of SN 2021lwz points to a very faint dwarf galaxy.
We model the observed spectral energy distribution (black data points in Figure 13) with the software package Prospector version 1.3 (Leja et al. 2017).151515Prospector uses the Flexible Stellar Population Synthesis (FSPS) code (Conroy et al. 2009) to generate the underlying physical model and python-fsps (Foreman-Mackey et al. 2014) to interface with FSPS in python. The FSPS code also accounts for the contribution from the diffuse gas based on the Cloudy models from Byler et al. (2017). We use the dynamic nested sampling package dynesty (Speagle 2020) to sample the posterior probability. We assume a Chabrier IMF (Chabrier 2003) and approximate the star formation history (SFH) by a linearly increasing SFH at early times followed by an exponential decline at late times [functional form , where is the age of the SFH episode and is the -folding timescale]. The model is attenuated with the Calzetti et al. (2000) model. The priors of the model parameters are set identically to those used by Schulze et al. (2021). Figure 13 shows the observed SED (black data points) and its best fit (grey curve). The SED is adequately described by a galaxy template with a mass of , and a star-formation rate of .
The mass and the star-formation rate are comparable to common star-forming galaxies of that stellar mass (grey band in Figure 14; Elbaz et al. 2007)161616The relationship between star-formation and galaxy mass shown in Figure 14 is based on results from Brinchmann et al. (2004) and is valid at ., and they are also similar to those of the host galaxy populations of SNe from the PTF survey (Schulze et al. 2018) albeit at the very low-mass end of the observed distribution. The data relative to USSNe compiled in Table 10 are given in Appendix J.
As a consistency check, we computed a plot similar to the one shown in Figure 13, without taking into account the emission lines. We find that the galaxy mass is almost unchanged. On the other hand the SFR departs from a factor 10 from the one discussed in the text and shown in Figure 14. The new reduced chi-square is also higher. All these points make us confident that, despite the low number of data points, the fit shown in Figure 13 including the emission-lines is better suited than a fit that would not include these lines.
5 Discussion
Based on our previous modelling and analysis it is not clear what is the nature of SN 2021lwz. This very fast evolving explosion (t days, rest-frame), with a low ejecta mass ( ), is drawn from a very low-brightness ( mag), low mass () dwarf galaxy of SFR () about ten times larger than the typical star-forming galaxies.
| Object | R.A. | Decl. | Type | z | trise,rf (a) | band | (b) | (c) | References |
| Name | [days] | [mag] | [mag] | ||||||
| SN 2018bgv | 11:02:30.29 | +55:35:55.79 | SLSN-I | 0.08 | g | 17.06 | -20.71 | 1 | |
| AT 2020xnd | 22:20:02.03 | -02:50:25.27 | LFBOT | 0.2442 | r | 19.56 | -20.72 | 2 | |
| SN 2014ft | 23:33:27.95 | +33:38:46.14 | USSN | 0.1868 | r | 19.62 | -20.06 | 3 | |
| iPTF 16asu | 12:59:09.28 | +13:48:09.19 | SL Ic-BL | 0.187 | g | 19.14 | -20.53 | 4, 5 | |
| SN 2018gep | 16:43:48.22 | +41:02:43.37 | SL Ic-BL | 0.031875 | 5.40.5 | g | 15.93 | -19.84 | 6 |
| SN 1998bw | 19:35:03.310 | -52:50:44.81 | Ic-BL | 0.008499 | 17 | V | 13.60 | -19.33 | 7, 8 |
| SN 2021lwz | 09:44:47.40 | +34:42:44.25 | ? | 0.065 | 7 | g | 17.25 | -20.08 | This work |
To discuss SN 2021lwz in a broader context, in Figure 15 we show the rest-frame light curves of the different types of transients listed in Table 3; namely SN 2018bgv (SLSN-I), SN 1998bw (peculiar Ic-BL), AT 2020xnd (LFBOT), SN 2014ft (USSN), iPTF16asu (SL Ic-BL) and SN 2018gep (SL Ic-BL). As it is done for SN 2021lwz, if the photometry is not already corrected for the extinction contribution from our Galaxy, it is done using Schlafly and Finkbeiner (2011). As much as possible the filters were chosen by taking into account the redshift of each source so that the rest-frame wavelength ranges covered overlap strongly with each other.
From the comparison of their rising time parameters given in Table 3 all these transients, except SN 1998bw, reach their maximum light in about 10 or less days. SN 2018bgv appears slightly more energetic than SN 2021lwz. On the other side of the absolute magnitude range, SN 1998bw is the less energetic transient. SN 2018bgv and SN 1998bw whose rest-frame rising times are about twice or more the rising time of SN 2021lwz, both have smooth light curves decreasing slightly faster than the one of SN 2021lwz. The light curves of iPTF 16asu, SN 2014ft, AT 2020xnd SN 2028gep, SN 2014ft and SN 2021lwz, all strongly overlap with each other before and up to a few days after maximum light. A few days after peak, the light curves of AT 2020xnd, SN 2014ft and SN 2018gep decrease faster than the light curves of SN 2021lwz, though. All in all, the ligth curve of iPTF 16asu is the one looking more similar to the one of SN 2021lwz.
5.1 H-poor SLSNe
SN 2018bgv is the fastest evolving ( days rest-frame) Type-W SLSN-I reported in the literature. It peaks close to mag. It is therefore very interesting for comparisons with SN 2021lwz. While we didn’t model it with our magnetar model, the light curve and a spectral sequence have been studied in detail by Lunnan et al. (2020). From black-body fits analysis of the light curve the authors report a photosphere temperature close to 11,000 K and a black-body radius close to 2 cm, a few days after peak, i.e. a photopshere with a radius about twice larger than, and a temperature about similar to, those of SN 2021lwz photosphere at similar phases. Close to peak the velocity of the photosphere of SN 2018bgv appears to be about 1.5 times the velocity estimated for SN 2021lwz. A fit of the light curve with the magnetar spin-down model using mosfit infers an ejecta of mass, , i.e. about 5 times higher than the one of SN 2021lwz as from our modelling estimate, hence suggesting the conditions behind the explosions of SN 2018bgv and SN 2021lwz are different.
As discussed in Section 3.1 SN 2018bgv provides a relatively good match with SN 2021lwz close to maximum light, but this is obtained at a lower redshift z than the redshift of SN 2021lwz. This is mainly because SNID tries to adjust the archetypal W-shape feature attributed to O ii, before and close to maximum light, observed in many Type I SLSNe. Indeed, and as discussed in Section 4.1, for H-poor SLSNe the typical absorption spectral features expected to be observed before maximum light, between 4000 and 5000 Å (see Könyves-Tóth and Seli 2023, Figure 1), in Type 15bn (e.g. Fe iii) and in Type W (e.g. O ii), are not detected in the pre-maximum spectra of SN 2021lwz. Instead, C ii, O i, Na i, Mg ii, Si ii, Ca ii, Ti ii and Fe ii are included to reproduce the shape of these spectra, and this despite the fact that from the estimates of the photosphere temperature K and the photosphere velocities kms before maximum light, one would expect features from a W-Type SLSN-I (see Könyves-Tóth 2022, Figure 7). We point out, however, that the conditions for these lines to appear are likely limited to a more stringent range of temperature K once spectral modelling effects, which are not included in the SYN++ modelling, are taken into account (see Saito et al. 2024). However, the study provided by Aamer et al. (2025), obtained on a sample larger than the one used by Könyves-Tóth and Seli (2023) suggests a continuum of spectral properties, hence that the differences of spectral profiles in type I SLSNe may not correspond to physically different type of events. In this larger sample Fe ii lines appear to be an important spectral marker formed near the photosphere as it is also found for SN 2021lw.
5.2 LFBOTs
The LFBOT AT 2020xnd and SN 2021lwz are peaking at about similar absolute magnitudes (see Figure 15), but AT 2020xnd is a faster evolving event than SN 2021lwz, mainly because it has a higher initial pulsar luminosity, assuming a magnetar model. Two spectra of AT 2020xnd at phases close to +5 days and +7 days after maximum light are publicly available in the WISeREP repository 181818URL to AT 2020xnd public spectra in WISeREP, https://www.wiserep.org/object/18199.. From the point of view of the photosphere evolution, the two very smooth spectra, which are continuum blue flux dominated do not show any strong absorption features, unlike those seen in the spectra of SN 2021lwz after maximum light (see Figure 3). This observational fact, combined with the analysis of their light curves using the magnetar model (for AT 2020xnd see Table 2 in Omand and Sarin 2024), suggests that AT 2020xnd and SN 2021lwz differ in nature, and do not pertain to the same class of transients. This conclusion is sustained by a visual comparison of some spectra of SN 2021lwz with some of SN 2020xnd and SN 2018cow as shown in Figure 16. The spectral evolution of SN 2021lwz show absorption features stronger than those observed in the spectra of AT 2020xnd and AT 2018cow which, in comparison, look almost featureless.
5.3 USSN
SN 2014ft has been observed, studied and classified by De et al. (2018) as a USSN that likely formed a compact neutron star binary. It is recognised as the first ever detection of a USSN (Hijikawa et al. 2019), while previous studies of fast evolving Type Ic-BL SNe like SN 2005ek (Drout et al. 2013) and SN 2010X (Kasliwal et al. 2010) pointed-out the existence of this class of objects. Posterior to these works, SN 2019dge (or ZTF18abfcmjw) was added to this class of transients as discussed by Yao et al. (2020). More recently SN 2023zaw has been shown to be a fast evolving USSN of very low ejecta mass ( ) (Das et al. 2024; Moore et al. 2025).
SN 2014ft is modelled using the Redback (Sarin et al. 2024) magnetar-driven supernova model in Omand and Sarin (2024). SN 2021lwz and SN 2014ft have very similar light curves both peaking at absolute magnitudes close to -20.1 mag (Figure 15), and similar fast rising times of order 6–7 days (see Table 3. Assuming the magnetar-driven supernova model is appropriate to describe both explosions, the posterior parameters from the light curve modelling both infer very low ejecta of the same order ( ), and supernovae explosion energies of the same order ( erg ). The main difference is that SN 2021lwz has an initial spin-down luminosity about 100 times higher than the one of SN 2014ft, and a spin-down timescale about 10 times faster. The braking index of both explosions being of the same order (), this results in SN 2021lwz having a total rotational energy an order of magnitude higher than SN 2014ft.
During the first 30 hours after explosion SN 2014ft showed He ii 4686 Å lines in 3 spectra (see Figure3 A in De et al. 2018). Such early spectra are not available for SN 2021lwz therefore we don’t know if a similar feature was present very early in its photosphere. A similar feature is not detected in the earliest spectra available on SN 2021lwz obtained about 5 to 7 days after the explosion (see Figure 3). Hints of He i, but not of He ii, are only suggested by the modelling analysis in the post peak spectra of SN 2021lwz. At maximum light the spectra of SN 2021lwz and SN 2014ft show common ionization atomic lines from Ca ii, Mg ii, Fe ii, Ti ii, Si ii, C ii, and O i (see our SYN++ analysis and Figure 5B in De et al. 2018) but absorption lines of P iii (while not unambiguous) and Na i observed in the spectrum of SN 2021lwz are not detected in the spectrum of SN 2014ft. The distinct temperatures of the photosphere estimated for each supernova close to maximum light, with K for SN 2021lwz (this work), and K for SN 2014ft (Figure 4B in De et al. 2018), may explain this difference. Before and near maximum light SN 2021lwz has a steeper continuum than SN 2014ft, showing that it is hotter and likely coming from a more energetic event, an inference consistent with the one provided by the modelling of the light curves. In the near-maximum spectra, around 5000 Å, SN 2014ft shows a ‘basin shaped’ feature similarly to SN 2015bn, while this is missing from the spectra of SN 2021lwz, here again most probably because of the difference in the temperatures. The post-maximum spectral lines are somewhat more similar, but SN 2014ft is still cooler than SN 2021lwz.
Finally, we point out that, contrary to the high specific SFR of the host of SN 2021lwz, the specific SFR of USSNe shown in Figure 14 is generally lower than, or about, the average expected value for typical star-forming galaxies (grey band). The stellar masses of the USSNe galaxies are also 2 to 4 orders of magnitudes higher than the mass of SN 2021lwz’s host. The statistics is low with only 4 hosts of USSNe, though.
All the arguments put together it is not clear if SN 2021lwz has been ultra-stripped or not, the main argument in favour of an ultra-stripped process being that it has a low ejecta mass, as does have SN 2014ft.
5.4 Typical H-poor stripped envelope SNe
Hydrogen-poor, stripped envelope, core-collapse supernovae (SESNe), of Types Ib, Ic and Ic-BL, are well studied in the optical (e.g. Drout et al. 2011; Taddia et al. 2019) the NIR (e.g. Shahbandeh et al. 2022), and at radio wavelengths (Srinivasaragavan et al. 2024). As has been discussed previously our post-peak optical spectral modelling hints at a marginal detection of He i, and the possible 1.0830 m line. Possible connections and a continuity between typical Type Ib and Ic has been discussed with no clear results about residual detections of He i in optical spectra (e.g. Hachinger et al. 2012; Piro and Morozova 2014; Prentice et al. 2018; Dessart et al. 2020; Williamson et al. 2021). More recently, Shahbandeh et al. (2022) presented the analysis of 75 NIR spectra of 34 SESNe with the aim to discuss NIR strong He i lines: m and m. In all spectra, except a few outliers, most of the absorption lines features observed close to 1 m can be explained by a combination of strong lines (C I m, He i m, and Mg II m), and the authors concluded that there is no distinct difference between He-poor and He-rich groups in this wavelength range. On the other hand, by analysing the 2 m features the authors showed one can clearly discriminate He-rich (optical IIb/Ib) from He-poor (optical Ic/Ic-BL) SESNe and computed SESNe templates using a Principal Component Analysis. The +13 days templates of such He-rich and He-poor SNe are shown in Figure 4 for illustration. A comparison with the +12.6 days NIR spectrum of SN 2021lwz, which optical spectral modelling points out to a Type Ic-like, disfavours the presence of residual He i. Actually, in the vicinity of the He i 2 m feature, the spectrum of SN 2021lwz looks flatter than the He-poor template, as looks the spectrum of the peculiar Type Ic-BL, SN 1998bw.
Additional arguments making SN 2021lwz peculiar with respect to SESNe are the following. It is more luminous at peak than the majority of these objects (see for example our plot in Figure 1 for a direct comparisons with Type Ic and Ic-BL absolute magnitudes). It has a low mass ejecta (as inferred from our multi-band modelling), , which is in contradiction with typical 56Ni ejecta masses of M and 0.5 for SNe Ib/c and SN Ic-BL, respectively (Drout et al. 2011), as well as with the ejecta masses expected for these two Types of SNe, of order M and 5 , respectively (Drout et al. 2011; Taddia et al. 2019). More recently Rodríguez et al. (2023) studied 191 stripped-envelope SNe, including 50 SNe Ib and 96 SNe Ic, of which 26 are SNe Ic-BL. Compared to that SN Ib/Ic/Ic-BL sample, SN 2021lwz is at least ten times more luminous at peak than the average (2.7 and 4.5 erg s-1 for Ib and Ic/Ic-BL, respectively). Its ejecta mass is about ten times lower than the average (2.6 ), and also lower than the minimum ejecta mass (0.5 ), making it a low occurrence event in the bottom tail of the distribution.
5.5 Comparisons with peculiar Type Ic-BL SN 1998bw
Before and close to maximum light the photosphere of SN 2021lwz looks more similar to a normal Type Ic supernova than to any other Type, with C ii, O i, Na i, Mg ii, Si ii, Ca ii, Ti ii and Fe ii lines identified in its earliest spectrum (see Section 4.1), and the velocity of the spectral features and of the photosphere are close to those of a Type Ic. Slightly broader optical spectral features observed after peak suggest that SN 2021lwz could have evolved closer to a Type Ic-BL. Actually, before maximum light, the SYN++ modelling shows that the photosphere velocity of SN 2021lwz is not extraordinarily high ( km/s), and the broadness parameter (see values displayed in Table 9) are pointing on a normal Type Ic. After maximum light, the input parameter value had to be higher to adjust the formation of several broader spectral features, making SN 2021lwz may be closer to the limit of a Ic-BL. But still, after peak, the best match is obtained with LSQ12dlf which is a SLSN-I showing velocity lines below the 30000 km/s seen in classic SNe Ic-BL. The match with the +14 days and +16 days spectra of SN 1998bw are not too bad above but are relatively poor at lower wavelengths. Some attempts to fit the -3.4 days Keck spectrum of SN 2021lwz with early spectra of SN 1998bw proved unsuccessful. Some differences between the NIR spectra of both transients can also seen in Figure 4. All in all, SN 2021lwz looks more like a SLSN showing the Ic-SL Type after peak.
5.6 Comparisons with fast evolving luminous Type Ic-BL iPTF16asu
iPTF 16asu whose origin is not yet well understood has been first discussed by Whitesides et al. (2017). It is a luminous, rapidly evolving, stripped-envelope SN peaking at mag. With a rest-frame rise time estimated as 4 days it appears as a faster evolving event than SN 2021lwz. Contrary to SN 2021lwz which looks like a Type Ic SN before peak, iPTF 16asu shows a relatively featureless blue continuum, but contrary to SN 2021lwz, iPTF 16asu shows Ic-BL features after peak (see some spectra shown in Figure 17 for comparison).
In Appendix I we present the fit of the multi-band light curve of iPTF 16asu using the same software, model, and priors as for SN 2021lwz (see Section 4.3 for details). The light curve fit is shown in Figure 22 and posterior is shown in Figure 23. From our modelling, the estimated SN explosion energy of iPTF 16asu is about 30 times stronger than the one of SN 2021lwz, which could explain the optical spectral differences before peak.
More in depth light curve modelling of iPTF 16asu has been discussed by Wang et al. (2019). In Section 4.2 we discuss the difficulty to model the light curve of SN 2021lwz using the Arnett model, similarly the analysis proposed by these authors underlined the difficulty to model the light curve of iPTF 16asu with radioactive heating (56Ni-only), and also the magnetar 56Ni models. Our modelling of iPTF 16asu and SN 2021lwz includes a fraction of Nickel in the ejecta mass (see Equation 9 in Omand and Sarin 2024), which given the total ejecta mass estimate and , of each object, respectively, means the Nickel mass would be excessively low. Energetically speaking, we find that these low ejecta masses are both associated to rotational energies of the same order (). The light curves of both events are very similar and peak at absolute magnitudes close to -20.3 mag. They also have fast rising times sharing a common parameter space ( days; see Table 3 and Figure 1). The g-band data plotted in Figure 15 show the steep increase of iPTF 16asu, while the r-band data show a decreasing light curve very similar to the one of SN 2021lwz.
The interpretation of the results obtained by Wang et al. (2019), and additional modelling of the light curves has been proposed by Wang and Gan (2022). The construction of the theoretical bolometric light curve is revised and the authors find that the magnetar plus 56Ni model can now account for the photometry of iPTF 16asu with a peak luminosity inferred from the theoretical light curve of erg. The best fit value for the ejecta mass is and the best fit parameter of the Nickel mass is . Some results very close to the best fit parameters obtained with our modelling of the multi-light curve of SN 2021lwz. All these results possibly make SN 2021lwz part of a class of rare rapidly evolving ( week), luminous type Ic-BL SNe ( mag or slightly higher), like iPTF 16asu (see Wang and Gan 2022), even though, SN 2021lwz is an intermediate velocity ( km s-1) type Ic/Ic-BL object, while iPTF 16asu is a high velocity type Ic-BL after peak.
5.7 Comparisons with fast evolving luminous Type Ic-BL SN 2018gep
SN 2018gep (see Ho et al. 2019; Leung et al. 2021; Pritchard et al. 2021) is another rapidly rising (1.4 0.1 mag hr-1 during the first hours just after explosion), stripped envelope SN, spectroscopically classified as a Ic-BL (see Ho et al. 2019). As a Ic-BL SN its light curve shows similarities with the one of iPTF 16asu (see Pritchard et al. 2021, and Figure 15). With a total fast rise time, days after explosion and a peak absolute magnitude mag it shares a common locus with iPTF 16asu and SN 2021lwz on the extreme side of the distribution of SN Ic-BL in this parameter space (see Pritchard et al. 2021, and our Figure 1). The spectra obtained during the first three days after explosion are blue dominated spectra showing smooth absorption features of highly ionized elements (CIV, O III, CIII) and a very hot photosphere decreasing from K to K. Then the W-Type signature attributed to O II and C II absorption lines observed in many Type I SLSNe before and around peak is visible in the spectra of SN 2018gep obtained during the fourth day after explosion (see Figure 11 in Ho et al. 2019). This marks the main difference with SN 2021lwz spectra in which this signature is interpreted differently given the assumed redshift of SN 2021lwz (see Section4.1).
All in all, all these comparisons suggest that SN 2021lwz has characteristics of, and is probably closely related to, many of these classes, and is therefore most likely some kind of transitional event.
5.8 Prospects and final remarks
An interesting test of the magnetar scenario for luminous, fast evolving objects is non-thermal emission. These objects could emit luminous X-ray (Kotera et al. 2013; Metzger et al. 2014; Murase et al. 2015) and radio emission (Murase et al. 2015; Omand et al. 2018; Eftekhari et al. 2021; Murase et al. 2021; Omand et al. 2025a), similar to PTF 10hgi (Eftekhari et al. 2019; Law et al. 2019; Mondal et al. 2020) or SN 2012au (Stroh et al. 2021) after a few years, but on timescales of a few weeks to a year. There could also be bright infrared emission on short timescales due to dust heated by the PWN (Omand et al. 2019), although this emission would likely fade quickly as the dust cools. The dominant non-thermal component can be predicted, in principle, from the nebular spectrum of the supernova, since line emission is sensitive to the underlying PWN spectrum (Omand and Jerkstrand 2023); however, these models are not currently capable of reproducing nebular spectra on timescales of less than a year.
The multi-dimensional structure of these magnetar-driven supernovae could show strong clumping due to Rayleigh-Taylor instabilities between the nebula and ejecta (Blondin and Chevalier 2017; Suzuki and Maeda 2017). When , such as in SN 2021lwz, the shock can break through the ejecta (Blondin and Chevalier 2017), leading to a decrease in the ejecta acceleration efficiency. This process, known as blowout, has been inferred to happen in the Crab Nebula (Omand et al. 2025b). Clumpy ejecta can cause polarization if the clump size is large enough (Tanaka et al. 2017), but it is currently unknown how to determine the clump size distribution from the inferred engine properties of the supernova.
It has been proposed by Zou and Cheng (2018) that Type I SLSNe and Type Ic-BL could be produced by jets. If so, they would both show strong polarization from early times. We point out that for SN 2021lwz our non-detections contribute to rule out this scenario. Additionally, if the energy was injected into the ejecta early, most of it would accelerate the ejecta instead of contributing to the luminosity (Suzuki and Maeda 2021).
6 Conclusions
SN 2021lwz is a peculiar luminous supernova at a redshift z=0.065 of absolute magnitude - 20.1 (AB system) discovered in a very low-brightness ( mag), low mass () dwarf galaxy of SFR () about ten times larger than the typical star-forming galaxies.
An extensive analysis and modelling of the photometry, spectroscopy and polarimetry data of the SN, followed by comparisons with SNe of well know types and with rarer transients, consistently leads to the following results:
SN 2021lwz shows a rapidly evolving and luminous optical light curve. The bolometric light curve rises in about days to a peak luminosity of about erg/s, at a rate of 0.2 mag day-1 close to the peak. From the intrinsic bolometric luminosity we derive days, and days, rest frame.
SN 2021lwz has a low mass ejecta (as from the magnetar model) with estimated velocities of km s-1 at days and of km s-1 after peak at and days, respectively. Near peak the photosphere has a temperature of K, a radius of cm and a velocity of 10 000 km s-1, consistent with the ejecta velocity.
If SN 2021lwz is a magnetar-driven supernova it is expected to have an initial luminosity erg s-1, a spin-down timescale day, and an ejecta mass, a final ejecta scale velocity, and a photospheric temperature at peak, of values similar to the ones reported above.
The magnetar-driven model leads to a very short diffusion time, 5 days, due to the low ejecta mass, to a bolometric peak timescale 7 days and to a -band peak timescale 10 days; the discrepancy between these values (Figure 12) is consistent with the high photospheric temperature and rapid cooling around peak (Figure 7). The SN is expected to have entered the nebular phase on the timescale of days but the spectroscopy shows the nebular phase should have started after +30 days. The gamma-ray leakage timescale, , is very long.
The shape of the photosphere of SN 2021lwz looks close to perfectly spherical and does not show any loss of symmetry between the two phases close to ( +1.0 days) and after maximum light ( +19. days). The observed symmetry of the system rules out a jet activity to be at the origin of SN 2021lwz.
Spectroscopically, SN 2021lwz looks like a Type Ic-like SN before maximum light. After maximum light some velocity lines may be a bit larger but the post-peak spectrum looks most like a SLSN and the continuum slope (see Figure 3) is much more like a SLSN that a Ic-BL, which is usually much redder from line blanketing.
Given its spectroscopic properties and the low ejecta mass needed to model its light-curve, SN 2021lwz does not match with many core-collapse H-poor SNe Types.
Using the magnetar model, the low ejecta mass of SN 2021lwz and its light curve shape possibly make it a USSN but it is not possible to demonstrate it with the data at hand.
SN 2021lwz shows similarities with the peculiar Type Ic-BL SN 1998bw but some differences in their rise time, absolute magnitudes at peak and in their NIR spectra suggest they are two different types of objetcs.
The light curve and multi-band light curve modelling of SN 2021lwz suggest it shares similarities with iPTF 16asu, and with SN 2018gep. The three sources show continuum blue dominated spectra before maximum light, even though SN 2021lwz shows spectral features pointing to a Type Ic-like SNe, while iPTF 16asu and SN 2018gep don’t. A possible explanation is that iPTF 16asu and SN 2018gep are more energetic explosions quickly rising to higher photosphere temperatures than is SN 2021lwz.
In conclusion SN 2021lwz is an uncommon transient showing many similarities with several classes of transients, and with rare transients. It may be an interesting example pointing on how differences in ejecta mass and engine parameters could produce a wide range of engine-driven SESNe.
With the advent of new high-cadence large-scale deeper surveys like the Vera C. Rubin observatory Large Synoptic Survey Telescope (LSST; Ivezić et al. 2019), very early detections of bright, low redshift events similar to SN 2021lwz may be helpful to better understand such phenomena. Similarly, data obtained with the Euclid satellite (Euclid Collaboration et al. 2025) should be helpful to tackle such problems. The observing strategy is not optimized for transient science but the Quick Data Release 1 has proven to be useful for early and late type detections of some SNe (see Duffy et al. 2025).
Acknowledgements.
The authors would like to thank the anonymous referee for her/his insightfull and constructive comments, Sheng Yang, Anna Ho and Ósmar Rodríguez for helpful discussions and inputs. We thank Rosa Clavero and David Morate for their kind support on the NOT. We thank George Khorenzhev (Space Research Institute, Russian Academy of Sciences) for contributing to some observations. F.P. and C.G.C. acknowledge support from the Spanish Ministerio de Ciencia, Innovación y Universidades (MICINN) under grant numbers PID2022-141915NB-C21. R.K.T. has been supported by the NKFIH/OTKA FK-134432 grant of the National Research, Development and Innovation Office of Hungary (NKFIH). C. M. B. O. acknowledges support from the Royal Society (grant Nos. DHF-R1-221175 and DHF-ERE-221005). N. Sarin acknowledges support from the Kavli foundation. T.-W.C. acknowledges the Yushan Fellow Program by the Ministry of Education, Taiwan for the financial support (MOE-111-YSFMS-0008-001-P1). T.K. acknowledges support from the Research Council of Finland project 360274. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute of Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, IN2P3, University of Warwick, Ruhr University Bochum, Cornell University, and Northwestern University. Operations are conducted by COO, IPAC, and UW. Zwicky Transient Facility access was supported by Northwestern University and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA). The ZTF forced-photometry service was funded under the Heising-Simons Foundation grant No. 12540303 (PI: Graham). SED Machine is based upon work supported by the National Science Foundation under Grant No. 1106171. The Gordon and Betty Moore Foundation, through both the Data-Driven Investigator Program and a dedicated grant, provided critical funding for SkyPortal. Based on observations made with the Nordic Optical Telescope (NOT), owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofísica de Canarias. The data presented here were obtained in part with ALFOSC, which is provided by the Instituto de Astrofísica de Andalucia (IAA) under a joint agreement with the University of Copenhagen and NOT. Some of the data were obtained during CAT service observation Spanish time. ALFOSC polarimetry imaging data of SN2020ank and calibration data were retrieved from the NOT public archive. The Liverpool Telescope is operated on the island of La Palma by Liverpool John Moores University in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias with financial support from the UK Science and Technology Facilities Council. This work has made use of data from the Asteroid Terrestrial- impact Last Alert System (ATLAS) project. The Asteroid Terrestrial- impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; by-products of the NEO search include images and catalogues from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. Lasair is supported by the UKRI Science and Technology Facilities Council and is a collaboration between the University of Edinburgh (grant ST/N002512/1) and Queen’s University Belfast (grant ST/N002520/1) within the LSST:UK Science Consortium. SNID is Copyright (C) 1999-2007 Stéphane Blondin and John L. Tonry, and is available under the GNU General Public License. This research made use of the Transient Name Server (TNS) which is the official IAU mechanism for reporting new astronomical transients such as supernova candidates, As of January 1, 2016. This research made use of Legacy Survey Data Release 10 data. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID #2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID #2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID #2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas AM University, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientifico e Tecnologico and the Ministerio da Ciencia, Tecnologia e Inovacao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenossische Technische Hochschule (ETH) Zurich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciencies de l’Espai (IEEC/CSIC), the Institut de Fisica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig Maximilians Universitat Munchen and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas AM University. The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract; and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO.Data Availability
For science reproducibility purposes, the photometry and the spectra presented in this work are available via WISEReP191919https://www.wiserep.org/object/18321. (Yaron and Gal-Yam 2012).
Appendix A Host Photometry Table
The measurements in the different bands of the host galaxy of SN 2021lwz discussed in Section2.1, are given in Table4. The magnitudes are not corrected for extinction.
| Survey | Band | Brightness |
| Telescope/Instrument | (mag) | |
| CFHT/MegaPrime | ||
| SDSS | ||
| CFHT/MegaPrime | ||
| Pan-STARRRS | ||
| SDSS |
Appendix B Spectroscopy Table
The log of the spectroscopy obtained on SN 2021lwz discussed in Section 2.2 is given in Table,5. The rest-frame phase relative to the g-band filter is given using MJD days.
| Instrument | Observation Date | Rest-frame phase |
| [UT] | (relative to g-band) | |
| P60/SEDM | 2021-05-11T04:07:47 | -5.4 |
| Keck/LRIS(∗) | 2021-05-13T05:40:20 | -3.4 |
| P60/SEDM | 2021-05-15T04:07:59 | -1.6 |
| NOT/ALFOSC | 2021-05-15T22:02:58 | -0.9 |
| Keck/LRIS | 2021-05-16T09:40:23 | -0.5 |
| NOT/ALFOSC | 2021-05-18T21:57:56 | 1.9 |
| P60/SEDM | 2021-05-24T04:00:59 | 6.8 |
| P60/SEDM | 2021-05-28T04:04:46 | 10.6 |
| P60/SEDM | 2021-05-30T05:48:20 | 12.5 |
| Keck/NIRES | 2021-05-30T07:15:08 | 12.6 |
| NOT/ALFOSC | 2021-05-30T21:58:39 | 13.2 |
| P60/SEDM | 2021-06-13T04:15:04 | 25.6 |
| NOT/ALFOSC | 2021-06-16T21:20:28 | 29.1 |
Appendix C Photometry Table
All the photometry obtained on SN 2021lwz is given in Table 6.
[…]
| (days) | (mag) | |||
| 59344.2 | -6.7 | g | 18.76 0.03 | ZTF |
| 59344.2 | -6.7 | r | 19.21 0.04 | ZTF |
| 59345.3 | -5.6 | o | 18.69 0.05 | ATLAS |
| 59346.2 | -4.7 | r | 18.35 0.02 | ZTF |
| 59346.3 | -4.6 | g | 18.01 0.04 | ZTF |
| 59347.0 | -3.9 | sdssg | 17.86 0.03 | LT |
| 59347.0 | -3.9 | sdssr | 18.22 0.04 | LT |
| 59347.0 | -3.9 | sdssi | 18.39 0.08 | LT |
| 59347.0 | -3.9 | sdssu | 17.85 0.16 | LT |
| 59347.0 | -3.9 | sdssz | 18.73 0.28 | LT |
| 59347.4 | -3.5 | UVW1 | 17.41 0.04 | Swift |
| 59347.4 | -3.5 | U | 17.31 0.05 | Swift |
| 59347.4 | -3.5 | B | 17.54 0.09 | Swift |
| 59347.4 | -3.5 | UVW2 | 17.89 0.04 | Swift |
[…]
| (days) | (mag) | |||
| 59347.4 | -3.5 | V | 17.97 0.18 | Swift |
| 59347.4 | -3.5 | UVM2 | 17.55 0.03 | Swift |
| 59347.9 | -3.0 | sdssg | 17.67 0.03 | LT |
| 59347.9 | -3.0 | sdssr | 18.01 0.03 | LT |
| 59347.9 | -3.0 | sdssi | 18.44 0.03 | LT |
| 59347.9 | -3.0 | sdssu | 17.37 0.04 | LT |
| 59347.9 | -3.0 | sdssz | 18.64 0.07 | LT |
| 59348.2 | -2.7 | r | 17.96 0.02 | ZTF |
| 59348.2 | -2.7 | g | 17.53 0.01 | ZTF |
| 59348.3 | -2.6 | c | 17.77 0.02 | ATLAS |
| 59348.3 | -2.6 | UVW1 | 17.34 0.04 | Swift |
| 59348.3 | -2.6 | U | 17.09 0.06 | Swift |
| 59348.3 | -2.6 | B | 17.38 0.09 | Swift |
| 59348.3 | -2.6 | UVW2 | 17.84 0.04 | Swift |
| 59348.3 | -2.6 | V | 17.91 0.23 | Swift |
| 59348.3 | -2.6 | UVM2 | 17.45 0.04 | Swift |
| 59348.9 | -2.0 | sdssg | 17.52 0.03 | LT |
| 59348.9 | -2.0 | sdssr | 17.86 0.03 | LT |
| 59348.9 | -2.0 | sdssi | 18.23 0.03 | LT |
| 59348.9 | -2.0 | sdssu | 17.16 0.01 | LT |
| 59348.9 | -2.0 | sdssz | 18.48 0.05 | LT |
| 59349.3 | -1.6 | UVW1 | 17.32 0.04 | Swift |
| 59349.3 | -1.6 | U | 17.15 0.06 | Swift |
| 59349.3 | -1.6 | B | 17.34 0.08 | Swift |
| 59349.3 | -1.6 | UVW2 | 17.9 0.04 | Swift |
| 59349.3 | -1.6 | V | 17.3 0.14 | Swift |
| 59349.3 | -1.6 | UVM2 | 17.52 0.04 | Swift |
| 59349.3 | -1.6 | o | 17.93 0.03 | ATLAS |
| 59349.9 | -1.0 | sdssg | 17.4 0.03 | LT |
| 59349.9 | -1.0 | sdssr | 17.73 0.04 | LT |
| 59349.9 | -1.0 | sdssi | 18.1 0.03 | LT |
| 59349.9 | -1.0 | sdssu | 17.1 0.05 | LT |
| 59349.9 | -1.0 | sdssz | 18.29 0.05 | LT |
| 59350.2 | -0.7 | g | 17.34 0.01 | ZTF |
| 59350.2 | -0.7 | r | 17.74 0.02 | ZTF |
| 59350.3 | -0.6 | UVW1 | 17.21 0.05 | Swift |
| 59350.3 | -0.6 | U | 17.01 0.06 | Swift |
| 59350.3 | -0.6 | B | 17.29 0.1 | Swift |
| 59350.3 | -0.6 | UVW2 | 17.86 0.05 | Swift |
| 59350.3 | -0.6 | V | 17.31 0.18 | Swift |
| 59350.3 | -0.6 | UVM2 | 17.55 0.06 | Swift |
| 59351.9 | 1.0 | sdssg | 17.27 0.03 | LT |
| 59351.9 | 1.0 | sdssr | 17.56 0.03 | LT |
| 59351.9 | 1.0 | sdssi | 17.86 0.05 | LT |
| 59351.9 | 1.0 | sdssu | 16.96 0.03 | LT |
| 59351.9 | 1.0 | sdssz | 18.12 0.09 | LT |
| 59352.3 | 1.4 | o | 17.65 0.05 | ATLAS |
| 59352.9 | 2.0 | sdssg | 17.31 0.07 | LT |
| 59352.9 | 2.0 | sdssr | 17.61 0.04 | LT |
| 59352.9 | 2.0 | sdssi | 17.86 0.08 | LT |
| 59352.9 | 2.0 | sdssu | 17.06 0.05 | LT |
| 59352.9 | 2.0 | sdssz | 18.2 0.07 | LT |
| 59353.9 | 3.0 | sdssg | 17.36 0.06 | LT |
| 59353.9 | 3.0 | sdssr | 17.6 0.04 | LT |
| 59353.9 | 3.0 | sdssi | 17.86 0.06 | LT |
| 59353.9 | 3.0 | sdssu | 17.2 0.15 | LT |
| 59353.9 | 3.0 | sdssz | 18.02 0.12 | LT |
| 59354.2 | 3.3 | i | 17.95 0.05 | ZTF |
| 59354.2 | 3.3 | r | 17.69 0.03 | ZTF |
| 59354.3 | 3.4 | o | 17.72 0.03 | ATLAS |
| 59354.4 | 3.5 | UVW1 | 17.58 0.06 | Swift |
| 59354.4 | 3.5 | U | 17.21 0.08 | Swift |
| 59354.4 | 3.5 | B | 17.08 0.1 | Swift |
| 59354.4 | 3.5 | UVW2 | 18.38 0.06 | Swift |
| 59354.4 | 3.5 | V | 17.28 0.19 | Swift |
[…]
| (days) | (mag) | |||
| 59354.4 | 3.5 | UVM2 | 17.94 0.05 | Swift |
| 59356.2 | 5.3 | g | 17.63 0.03 | ZTF |
| 59356.2 | 5.3 | r | 17.82 0.04 | ZTF |
| 59356.3 | 5.4 | o | 17.86 0.04 | ATLAS |
| 59356.8 | 5.9 | UVW1 | 18.11 0.07 | Swift |
| 59356.8 | 5.9 | U | 17.6 0.09 | Swift |
| 59356.8 | 5.9 | B | 17.6 0.14 | Swift |
| 59356.8 | 5.9 | UVW2 | 19.08 0.08 | Swift |
| 59356.8 | 5.9 | V | 17.33 0.2 | Swift |
| 59356.8 | 5.9 | UVM2 | 18.67 0.06 | Swift |
| 59357.0 | 6.1 | sdssg | 17.72 0.04 | LT |
| 59357.0 | 6.1 | sdssr | 17.84 0.04 | LT |
| 59357.0 | 6.1 | sdssi | 18.07 0.04 | LT |
| 59357.0 | 6.1 | sdssz | 18.13 0.07 | LT |
| 59357.9 | 7.0 | sdssg | 17.77 0.07 | LT |
| 59357.9 | 7.0 | sdssr | 17.95 0.03 | LT |
| 59357.9 | 7.0 | sdssi | 18.1 0.03 | LT |
| 59357.9 | 7.0 | sdssu | 17.66 0.04 | LT |
| 59357.9 | 7.0 | sdssz | 18.2 0.05 | LT |
| 59358.2 | 7.3 | i | 18.08 0.04 | ZTF |
| 59358.2 | 7.3 | r | 17.93 0.05 | ZTF |
| 59358.2 | 7.3 | g | 17.82 0.04 | ZTF |
| 59359.9 | 9.0 | sdssg | 17.92 0.09 | LT |
| 59359.9 | 9.0 | sdssr | 18.02 0.04 | LT |
| 59359.9 | 9.0 | sdssi | 18.17 0.03 | LT |
| 59359.9 | 9.0 | sdssu | 18.27 0.01 | LT |
| 59359.9 | 9.0 | sdssz | 18.32 0.06 | LT |
| 59360.3 | 9.4 | o | 18.04 0.06 | ATLAS |
| 59361.2 | 10.3 | g | 18.01 0.05 | ZTF |
| 59362.9 | 12.0 | sdssg | 18.07 0.02 | LT |
| 59362.9 | 12.0 | sdssr | 18.15 0.03 | LT |
| 59362.9 | 12.0 | sdssi | 18.32 0.03 | LT |
| 59362.9 | 12.0 | sdssu | 18.38 0.08 | LT |
| 59362.9 | 12.0 | sdssz | 18.37 0.05 | LT |
| 59363.3 | 12.4 | o | 18.1 0.04 | ATLAS |
| 59364.1 | 13.2 | U | 18.7 0.24 | Swift |
| 59364.1 | 13.2 | B | 17.79 0.21 | Swift |
| 59364.1 | 13.2 | UVW2 | 20.42 0.25 | Swift |
| 59364.1 | 13.2 | UVM2 | 20.2 0.2 | Swift |
| 59365.9 | 15.0 | sdssg | 18.2 0.03 | LT |
| 59365.9 | 15.0 | sdssr | 18.26 0.04 | LT |
| 59365.9 | 15.0 | sdssi | 18.4 0.04 | LT |
| 59365.9 | 15.0 | sdssu | 18.76 0.09 | LT |
| 59365.9 | 15.0 | sdssz | 18.46 0.04 | LT |
| 59366.9 | 16.0 | UVW1 | 19.61 0.13 | Swift |
| 59366.9 | 16.0 | U | 18.72 0.13 | Swift |
| 59366.9 | 16.0 | B | 18.56 0.2 | Swift |
| 59366.9 | 16.0 | UVW2 | 20.94 0.28 | Swift |
| 59367.3 | 16.4 | o | 18.06 0.05 | ATLAS |
| 59368.5 | 17.6 | UVW1 | 19.84 0.14 | Swift |
| 59368.5 | 17.6 | U | 18.72 0.13 | Swift |
| 59368.5 | 17.6 | B | 18.22 0.15 | Swift |
| 59368.5 | 17.6 | UVW2 | 20.61 0.14 | Swift |
| 59368.5 | 17.6 | V | 18.21 0.28 | Swift |
| 59368.5 | 17.6 | UVM2 | 20.53 0.12 | Swift |
| 59370.9 | 20.0 | UVW1 | 19.94 0.2 | Swift |
| 59370.9 | 20.0 | U | 18.73 0.17 | Swift |
| 59370.9 | 20.0 | B | 18.31 0.21 | Swift |
| 59370.9 | 20.0 | UVW2 | 20.85 0.22 | Swift |
| 59370.9 | 20.0 | V | 17.79 0.28 | Swift |
| 59370.9 | 20.0 | sdssg | 18.28 0.03 | LT |
| 59370.9 | 20.0 | UVM2 | 20.5 0.17 | Swift |
| 59370.9 | 20.0 | sdssr | 18.32 0.03 | LT |
| 59370.9 | 20.0 | sdssi | 18.45 0.03 | LT |
[…]
| (days) | (mag) | |||
| 59370.9 | 20.0 | sdssz | 18.5 0.07 | LT |
| 59371.3 | 20.4 | o | 18.41 0.06 | ATLAS |
| 59376.9 | 26.0 | sdssr | 18.68 0.04 | LT |
| 59376.9 | 26.0 | sdssi | 18.71 0.05 | LT |
| 59376.9 | 26.0 | sdssz | 18.62 0.08 | LT |
| 59377.9 | 27.0 | sdssg | 18.75 0.04 | LT |
| 59377.9 | 27.0 | sdssr | 18.74 0.04 | LT |
| 59377.9 | 27.0 | sdssi | 18.83 0.05 | LT |
| 59377.9 | 27.0 | sdssz | 18.73 0.07 | LT |
| 59381.9 | 31.0 | sdssg | 19.05 0.07 | LT |
| 59381.9 | 31.0 | sdssr | 18.97 0.05 | LT |
| 59381.9 | 31.0 | sdssi | 19.26 0.12 | LT |
| 59381.9 | 31.0 | sdssz | 19.04 0.14 | LT |
| 59390.2 | 39.3 | r | 19.8 0.14 | ZTF |
| 59391.2 | 40.3 | r | 19.79 0.14 | ZTF |
| 59394.2 | 43.3 | r | 19.89 0.15 | ZTF |
| 59396.2 | 45.3 | r | 19.86 0.11 | ZTF |
| 59397.2 | 46.3 | r | 20.12 0.21 | ZTF |
Appendix D Linear Polarimetry material
Linear polarimetry of SN 2021lwz obtained with ALFOSC in one of the V-band is shown in Figure 18. The log of the polarimetry obtained on SN 2021lwz using ALFOSC on the NOT (see Section 2.4) is presented in Table,7.
| UT Time | Object | Exp. Time | Filter | Seeing |
| [s] | [] | |||
| 2021-05-17 21:37:36 | 2021lwz | V | 1.9 | |
| 2021-05-17 22:53:56 | HD127769 | V | 1.5 | |
| 2021-05-17 22:57:30 | GD 319 | V | 1.3 | |
| 2021-06-04 21:05:14 | 2021lwz | V | 0.6 | |
| 2021-06-04 21:26:07 | GD 319 | V | 0.5 | |
| 2021-06-04 21:30:53 | HD127769 | V | 0.5 |
| Date | Source | ||||||||
| 2021-05-17 | BD52913 | 0.03 | -0.10 | 0.100.09 | … | … | … | … | … |
| – | HD127769 | 0.45 | -1.41 | 1.480.06 | 1.380.11 | 143.82.4 | … | … | … |
| – | STAR 1 | -0.06 | -0.23 | 0.230.15 | 0.160.18 | 117.432.4 | 0.160.18 | 26.232.4 | … |
| – | 2021lwz (f) | 0.01 | 0.14 | 0.150.10 | 0.240.14 | 47.415.9 | 0.240.14 | 136.215.9 | 0.090.14 |
| – | … | … | … | … | … | … | … | … | 0.00.14 (bc) |
| 2021-06-04 | BD52913 | -0.05 | -0.09 | 0.110.15 | … | … | … | … | … |
| – | HD127769 | 0.52 | -2.02 | 2.090.31 | 2.020.34 | 143.34.8 | … | … | … |
| – | STAR 1 | -0.33 | -0.20 | 0.390.21 | 0.300.26 | 100.524.9 | 0.300.26 | 9.924.9 | … |
| – | 2021lwz (f) | -0.14 | 0.20 | 0.240.11 | 0.300.18 | 53.417.2 | 0.300.18 | 142.817.2 | 0.430.18 |
| – | … | … | … | … | … | … | … | … | 0.250.18(bc) |
Appendix E Fits obtained with SNID
Appendix F Spectrum modelling with SYN++
The best-fit global parameter values of the SYN++ modelling of SN 2021lwz (see Section 4.1 )are given in Table,1. The best-fit local SYN++ parameters of the modelled spectra of SN 2021lwz are given in this appendix in Table,9.
| Ions | He i | C ii | O i | Na i | Mg ii | Si ii | Ca ii | Ti ii | Fe ii |
| -3.4 days phase | |||||||||
| -1.0 | -1.0 | 0.9 | -0.5 | -0.4 | -0.3 | -1.6 | -0.5 | ||
| 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | ||
| 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | ||
| aux | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | |
| 20.0 | 20.0 | 5.0 | 20.0 | 20.0 | 20.0 | 7.0 | 12.0 | ||
| +13.2 days and +29.1 days phase | |||||||||
| -0.3 | -0.8 | -0.3 | 0.3 | 0.0 | 0.0 | 0.1 | |||
| 7.5 | 7.5 | 7.5 | 7.5 | 7.5 | 7.5 | 7.5 | |||
| 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | |||
| aux | 10.0 | 5.0 | 3.0 | 1.0 | 2.0 | 5.0 | 3.0 | ||
| 8.0 | 20.0 | 8.0 | 8.0 | 20.0 | 15.0 | 8.0 | |||
Appendix G Bolometric Light curve Arnett modelling
In this section we give the details about the bolometric light curve modelling using the Arnett model (see Arnett 1982; Kasen and Bildsten 2010; Woosley 2010; West et al. 2023) which main conclusion (the light curve can’t be fit with ) is given in Section 4.2.
Using the lmfit.minimize least squares method we retrieve the following parameter values for the radioactive source model: M M⊙, M M⊙, v km s-1 (see Section 3.6), E erg, t days and the diffusion timescale was days. The resulting best fit model is shown in red in Figure 20. Since the total ejecta mass includes the mass of 56Ni these parameter results must be considered to be unphysical.
By re-parameterising the 56Ni mass as a fraction of the ejecta mass we then force the fit to be physical. However, as shown by the dashed red line in Figure 20, the fit converges poorly.
We also fit the bolometric light curve with the magnetar model. Parameter values are found to be M M⊙, B Gauss, P, v km s-1 (see Section 3.6), t days, the diffusion timescale days, and the spindown timescale days. The resulting best fit parameter model is plotted in blue together with the bolometric light curve in green in Figure 20.
Appendix H Magnetar-driven model of SN 2021lwz
We fit the mulit-band light curve of SN 2021lwz shown in Figure 11 using the software, model, and priors discussed in Section 4.3. The posterior is shown in Figure 23.
Appendix I Magnetar-driven model of iPTF16asu
We fit the mulit-band light curve of iPTF16asu using the same software, model, and priors as for SN 2021lwz (see Section 4.3 for details). The light curve fit is shown in Figure 22 and posterior is shown in Figure 23.
Appendix J USSNe Table
The list of USSNe objects displayed in Section 4.4 and their associated properties are given in Table 10.
| OBJECT | MASSMED | MASSSUP | MASSINF | SFRMED | SFRSUP | SFRINF | REF. | TYPE |
| SN2021agco | 9.41 | 9.42 | 9.41 | -0.70 | -0.68 | -0.72 | 1 | USSN |
| SN2019dge | 8.4 | 8.5 | 8.30 | -1.52 | -1.46 | -1.60 | 2 | USSN |
| SN2019wxt | 10.58 | 10.71 | 10.22 | 0.44 | 0.79 | 0.30 | 3 | USSN |
| SN2014ft | 10.48 | 10.53 | 10.43 | 0.69 | 0.88 | 0.36 | 4 | USSN |